# Foundations of Adaptive High-Level Tight Control of Prostate Cancer: A Path from From Terminal Disease to Chronic Condition

### v2

@@ IV.1 Clinical Data and Nonlinear Mixed Effect
We reanalyze the clinical data from [22], which includes $N=32$ patients undergoing Abiraterone treatment. Note that these patients have already received treatment to block male hormone production from the pituitary gland, which accounts for approximately $\sim 90\%$ of total production. Abiraterone is used as an additional therapy, to inhibit androgen production from other sources, including the adrenal glands and the tumor itself, which contribute the remaining $\sim 10\%$ of total production. Considerations in our analysis differ from those of [22] in certain direction:
Figure 1: Time series of PSA levels for all $N=32$ patients, for alternative normalizations. (A) The PSA level in ng/L. (B) The PSA level as normalized in Zhang et al. [22], in which the maximum value of any given PSA time series after the normalization is always at $1$. (C) The normalized PSA level is adjusted so that the value at the start of the first treatment in any given PSA time series is always set to $1$ after normalization.
  • In Fig. 1A we show the raw time series data for the PSA level (in ng/L) of all patients. It is very important to note that these data are widely different in scale (the maximum value in each of these PSA time series ranges from $\sim 2$ng/L to $\sim 200$ng/L), so before fitting the model to all patients, we need to normalize them to ensure comparability between patients. If not, the fit will then be overly controlled by the highly volatile patients. Here, we normalize each patient’s PSA measurement data using the PSA value at the time Abiraterone was first administered (see Fig. 1C), instead of the maximum PSA value as in [22] (see Fig. 1B). This choice for normalization not only still ensures comparability between PSA time series across all patients, but also allows for data updates without necessitating changes to previous entries.
  • We decide to relax the assumption in [22] that the cancer growth rates $r_{S}$ and $r_{R}$ are identical for all patients. This is consistent with the values estimated for patients receiving continuous drug treatment, ranging from $2\times 10^{-3}$ to $2\times 10^{-2}$/day [22]. Instead, we permit $r_{R}$ to vary among patients while treating the ratio of these rates, $r_{S}/r_{R}$, as universal. We do not pre-set $r_{S}/r_{R}$.
@@
  • The model is nonlinear in both common as well as person specific (mixed) factors. For this reason, we utilize an established statistical approach, the empirical Bayes approach of Lindstrom and Bates [28], to fit the model to the normalized PSA data within the framework of nonlinear mixed effects (NLME). The estimated parameters for fitting the model to the data are selected so as to maximize the observation likelihood, meaning they provide the best possible explanation for the patterns seen in the data. Not only do we have estimates for the two common parameters $\{\alpha_{RS},r_{S}/r_{R}\}$, we also obtain probability distributions for four patient-specific parameters
@@
  • We do not assign weights differently (and rather arbitrarily) to different events (e.g. the first data points after a treatment change, from off-drug to on-drug and vice versa, are weighted five times more than the others in [22]). Here, all data points are treated the same.
@@
To be more precise, let us denote the two common parameters as $\Theta\equiv\{\alpha_{RS},r_{S}/r_{R}\}$ and the four patient-specific parameters (besides $\tilde{\lambda}$) as $\theta\equiv\{\gamma,r_{R},x_{S}(0),x_{R}(0)\}$, and use the subscript $\mu$ to specify the individual patient, i.e. $\theta_{\mu}$. For each patient $\mu$ in [22], we know their given Abiraterone treatment schedule $\Lambda_{\mu}(t)$, and the measurement times $t_{m_{\mu}}$ (where $m_{\mu}$ labels the time point) at which they have their PSA level recorded, as well as the measured values $\text{PSA}_{\mu}(t_{m_{\mu}})$. If $m_{\mu}^{\text{(treat)}}$ is the time point when Abiraterone was first administered, then the normalized PSA level can be calculated with:
$\overline{\text{PSA}}_{\mu}(t_{m_{\mu}})=\text{PSA}_{\mu}(t_{m_{\mu}})/\text{PSA}_{\mu}(t_{m^{\text{(treat)}}_{\mu}})\ .$ (10)
For a given $\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}$, we can use Eq. (4) to obtain theoretically what trajectory the total cancer cell population should be, [^1] We employ the fourth-order Runge-Kutta method for numerical integration to derive the trajectory in MatLab R2023a [29]. which we denote as $x^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)$. The theoretical trajectory for the normalized PSA level, therefore, follows from Eq. (4) if $\tilde{\lambda}_{\mu}$ is known:
$\overline{\text{PSA}}^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)=\tilde{\lambda}_{\mu}x^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)/\text{PSA}_{\mu}(t_{m^{\text{(treat)}}_{\mu}})\ .$ (11)
@@ IV.2 Statistical Assumptions
To use the Lindstrom-Bates empirical Bayes (LBEB) approach for NLME [28], we need to make some statistical assumptions:
@@
    Here $A$, $B$ are two unknown positive constants, and $\mathcal{N}(0,1)$ represents the Gaussian distribution centered at 00 with a variance of $1$.
  • The four patient-specific parameters $\{r_{R},(1-\gamma),x_{S}(0),x_{R}(0)\}$ are assumed to be sampled from four log-normal distributions [30]. Note that we consider $(1-\gamma)$ to be sampled from a log-normal distribution, not $\gamma$, because observations (PSA levels before and after Abiraterone administration) suggest that $\gamma$ is typically much closer to $1$ than 00.
We then apply the LBEB approach [^2] We use the standard package nlmefit in MatLab R2023a [29] to algorithmically execute LBEB approach for NLME. to estimate the common parameters $\{r_{S}/r_{R},\alpha_{RS}\}$, the log-normal distributions of $\{r_{R},(1-\gamma),x_{S}(0),x_{R}(0)\}$, and the values $\{r_{R},\gamma,x_{S}(0),x_{R}(0)\}$ of each patient, and also the two constants $A$, $B$ that characterized our noise function, so that the observations across patients can be the most likely ones (maximum likelihood).
@@ IV.3 Best-Fit Results
We start by acknowledging that the clinical data is extremely noisy, which should be expected. The fit found in Zhang et al. [22], while not quantitative, does capture nearly all the qualitative behavior of every patients regarding PSA response on- and off-drug. However, the total SSE for all patients of this fit is very large, which motivates us to reanalyze their data. While we still consider two common parameters, we now use four patient-specific parameters instead of only two, resulting in a significantly better fit, as shown in Fig. 2 and Fig. 3. The total SSE has been reduced by more than a factor of three, from $170.68$ for Zhang et al. [22] fit to $57.52$ for our LBEB approach. The SSE for individual patients is also typically smaller using our approach (26 out of 32).
@@
Figure 2: Time series of PSA levels for the first 16 patients.. Each plot title shows the patient’s name, SSE for Zhang et al. [22] fit, and the LBEB fit. The cyan line marks drug use periods, orange dots represent normalized PSA Eq. (10), the light-blue dashed line shows Zhang et al. [22] best fit, and the blue line shows the LBEB best fit.
@@
Figure 3: Time series of PSA levels for the next 16 patients. Each plot title shows the patient’s name, SSE for Zhang et al. [22] fit, and the LBEB fit. The cyan line marks drug use periods, orange dots represent normalized PSA Eq. (10), the light-blue dashed line shows Zhang et al. [22] best fit, and the blue line shows the LBEB best fit.
@@
Figure 4: The relations between patient-specific parameters and the PSA level at the first drug treatment. Here we show the identified resistant cell growth rate $r_{R}$, drug effectiveness $\gamma$, and initial population values $\{x_{S}(0),x_{R}(0)\}$, and the PSA level when the total cancel population reaches its carrying capacity $\text{PSA}_{K}=\tilde{\lambda}$ for all patients. (A) The resistant cell growth rate $r_{R}$ and the drug ineffectiveness $1-\gamma$. The green continuous line shows a slight positive trend between these parameters in log-log plot ($R^{2}=0.10$ [31]), indicating a power-law relation $1-\gamma\propto r_{R}^{0.5\pm 0.5}$. (B) The initial values $\{x_{S}(0),x_{R}(0)\}$. The yellow dash line divides this two-dimensional parameter space into two: sensitive cell dominated (above) and resistant cell dominated (below), based on their influences on the growth of resistant cells at $t=0$. (C) The PSA levels at first drug treatment $\text{PSA}(t_{m^{\text{treat}}})$ and at carrying capacity $\text{PSA}_{K}$. The green continuous line shows a linear regression (that passes through the origin) between these parameters ($R^{2}=0.59$ [31]), indicating a linear relation $\text{PSA}(t_{m^{\text{treat}}})=\left(0.25\pm 0.05\right)\text{PSA}_{K}$. The yellow dash line divides this two-dimensional space into two: $\text{PSA}(t_{m^{\text{treat}}})/\text{PSA}_{K}>1/\alpha_{RS}$ (above) and $\text{PSA}(t_{m^{\text{treat}}})/\text{PSA}_{K}<1/\alpha_{RS}$ (below).
@@
We find the best common parameter values to be $r_{S}/r_{R}=2.97\pm 0.02$ and $\alpha_{RS}=5.2\pm 0.9$. This provides strong evidence that $r_{S}>r_{R}$. The large value for the competition coefficient $\alpha_{RS}$ between resistant and sensitive cells agrees with the findings of [22], indicating the robustness of this estimation even with different approaches. The patient-specific parameters are sampled from
@@ IV.4 Comparison between Models
For linear regression, the standard metric for “goodness-of-fit” is given by the $R^{2}$-value [31]. A direct generalization of this dimensionless metric for nonlinear models [32, 33] is commonly referred to as the Efron’s pseudo-$R^{2}$-value [34], is given by $\tilde{R}^{2}=1-\tilde{A}$. The coefficient of alienation $\tilde{A}$, defined as [33]:
@@
where $\text{SS}_{\text{data}}$ is the data variance (i.e. SSE with respect to the mean value). This measures the proportion of residual variation in the nonlinear model compared to a null model with no explanatory power, indicating how much the model deviates from an ideal fit. An ideal fit corresponds to $\tilde{R}^{2}=1$. While there are many other different proposals for goodness-of-fit depending on the specific characteristics of the nonlinear model being used [35, 36, 37], the Efron’s pseudo-$\tilde{R^{2}}$-value remains a useful metric due to its simplicity and ease of interpretation [33].
Since the number of parameters in the fit models that we use are not the same, i.e. $n_{params}=(2\times N)+2=66$ in Zhang et al. [22] and $n_{params}=(4\times N)+2=130$ in our approach, we need to calculate the adjusted Efron’s psuedo-$R^{2}$-value [38] instead:
@@
We have $n_{\text{data}}=678$ and $\text{SS}_{\text{data}}=176.16$. The closer $\tilde{R}^{2}_{adj}$ is to $1$, the more we can trust a fitted model. We obtain $\tilde{R}^{2}_{adj}\approx-0.08$ for Zhang et al. [22] fit and $\tilde{R}^{2}_{adj}\approx 0.60$ for our LBEB fit, which corroborates our approach.
@@ V Optimization of Treatment with Abiraterone
Within our adaptive therapy policy framework, our goal is to determine the optimal policy parameters $(\eta_{\uparrow},\eta_{\downarrow})$ that maximize the total achievable time $\tau_{\text{tot}}$ we can keep the cancer progression (with treatment) below the biomarker level $\text{PSA}_{\text{thr}}$, the level at which the patient begins to show symptoms, in order to avoid further complications. Consider the patients with two common parameters $\Theta=\{r_{S}/r_{R},\alpha_{RS}\}$ and the five patient-specific parameters $\{\text{PSA}_{K},\theta\}=\{\tilde{\lambda},r_{R},\gamma,x_{S}(0),x_{R}(0)\}$, as found according to Section IV.3. Note that $\text{PSA}_{K}=\tilde{\lambda}$ follows, as explained in Eq. (6). For simplification, we further assume that the PSA level is measured continuously, and that there is no noise in these measurements, so that the decision to turn on and off drug treatment based on the policy parameters as described in Eq. (8) is exact. Hence, we can calculate the total achievable time $\tau_{\text{tot}}$ below $\text{PSA}_{\text{thr}}$ using policy parameters $(\eta_{\uparrow},\eta_{\downarrow})$ for a patient $\mu$ with parameters $\{{\Theta,\text{PSA}_{K}}_{\mu},\theta_{\mu}\}$, [^3] We utilize the standard integrator ode23 in MatLab R2023a [29] to simulate Eq. (4) and implement the terminal condition, obtaining the total achievable time $\tau_{\text{tot}}$ at when the simulated PSA level crosses the threshold value $\text{PSA}_{\text{thr}}$. and we denote its full functional form as $\tau_{\text{tot}}^{\{{\Theta,\text{PSA}_{K}}_{\mu},\theta_{\mu}\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$. We can simplify this expression to $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$, and further drop all superscripts when it is clear which patient and what threshold are referred to.

### v3

@@ IV.1 Clinical Data and Nonlinear Mixed Effect
We reanalyze the clinical data from [zhang2022evolution], which includes $N=32$ patients undergoing Abiraterone treatment. Note that these patients have already received treatment to block male hormone production from the pituitary gland, which accounts for approximately $\sim 90\%$ of total production. Abiraterone is used as an additional therapy, to inhibit androgen production from other sources, including the adrenal glands and the tumor itself, which contribute the remaining $\sim 10\%$ of total production. Considerations in our analysis differ from those of [zhang2022evolution] in certain direction:
Figure 1: Time series of PSA levels for all $N=32$ patients, for alternative normalizations. (A) The PSA level in ng/L. (B) The PSA level as normalized in Zhang et al. [zhang2022evolution], in which the maximum value of any given PSA time series after the normalization is always at $1$. (C) The normalized PSA level is adjusted so that the value at the start of the first treatment in any given PSA time series is always set to 11 after normalization.
  • In Fig. 1A we show the raw time series data for the PSA level (in ng/L) of all patients. It is very important to note that these data are widely different in scale (the maximum value in each of these PSA time series ranges from $\sim 2$ng/L to $\sim 200$ng/L), so before fitting the model to all patients, we need to normalize them to ensure comparability between patients. If not, the fit will then be overly controlled by the highly volatile patients. Here, we normalize each patient’s PSA measurement data using the PSA value at the time Abiraterone was first administered (see Fig. 1C), instead of the maximum PSA value as in [zhang2022evolution] (see Fig. 1B). This choice for normalization not only still ensures comparability between PSA time series across all patients, but also allows for data updates without necessitating changes to previous entries.
  • We decide to relax the assumption in [zhang2022evolution] that the cancer growth rates $r_{S}$ and $r_{R}$ are identical for all patients. This is consistent with the values estimated for patients receiving continuous drug treatment, ranging from $2\times 10^{-3}$ to $2\times 10^{-2}$/day [zhang2022evolution]. Instead, we permit $r_{R}$ to vary among patients while treating the ratio of these rates, $r_{S}/r_{R}$, as universal. We do not pre-set $r_{S}/r_{R}$.
@@
  • The model is nonlinear in both common as well as person specific (mixed) factors. For this reason, we utilize an established statistical approach, the empirical Bayes approach of Lindstrom and Bates [lindstrom1990nonlinear], to fit the model to the normalized PSA data within the framework of nonlinear mixed effects (NLME). The estimated parameters for fitting the model to the data are selected so as to maximize the observation likelihood, meaning they provide the best possible explanation for the patterns seen in the data. Not only do we have estimates for the two common parameters $\{\alpha_{RS},r_{S}/r_{R}\}$, we also obtain probability distributions for four patient-specific parameters
@@
  • We do not assign weights differently (and rather arbitrarily) to different events (e.g. the first data points after a treatment change, from off-drug to on-drug and vice versa, are weighted five times more than the others in [zhang2022evolution]). Here, all data points are treated the same.
@@
To be more precise, let us denote the two common parameters as $\Theta\equiv\{\alpha_{RS},r_{S}/r_{R}\}$ and the four patient-specific parameters (besides $\tilde{\lambda}$) as $\theta\equiv\{\gamma,r_{R},x_{S}(0),x_{R}(0)\}$, and use the subscript $\mu$ to specify the individual patient, i.e. $\theta_{\mu}$. For each patient $\mu$ in [zhang2022evolution], we know their given Abiraterone treatment schedule $\Lambda_{\mu}(t)$, and the measurement times $t_{m_{\mu}}$ (where $m_{\mu}$ labels the time point) at which they have their PSA level recorded, as well as the measured values $\text{PSA}_{\mu}(t_{m_{\mu}})$. If $m_{\mu}^{\text{(treat)}}$ is the time point when Abiraterone was first administered, then the normalized PSA level can be calculated with:
$\overline{\text{PSA}}_{\mu}(t_{m_{\mu}})=\text{PSA}_{\mu}(t_{m_{\mu}})/\text{PSA}_{\mu}(t_{m^{\text{(treat)}}_{\mu}})\$. (10)
For a given $\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}$, we can use Eq. (4) to obtain theoretically what trajectory the total cancer cell population should be, [^2] We employ the fourth-order Runge-Kutta method for numerical integration to derive the trajectory in MatLab R2023a [MATLAB]. which we denote as $x^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)$. The theoretical trajectory for the normalized PSA level, therefore, follows from Eq. (4) if $\tilde{\lambda}_{\mu}$ is known:
$\overline{\text{PSA}}^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)=\tilde{\lambda}_{\mu}x^{\left\{\Theta,\theta_{\mu};\Lambda_{\mu}(t)\right\}}(t)/\text{PSA}_{\mu}(t_{m^{\text{(treat)}}_{\mu}})\$. (11)
@@ IV.2 Statistical Assumptions
To use the Lindstrom-Bates empirical Bayes (LBEB) approach for NLME [lindstrom1990nonlinear], we need to make some statistical assumptions:
@@
    Here $A$, $B$ are two unknown positive constants, and $\mathcal{N}(0,1)$ represents the Gaussian distribution centered at 0 with a variance of $1$.
  • The four patient-specific parameters $\{r_{R},(1-\gamma),x_{S}(0),x_{R}(0)\}$ are assumed to be sampled from four log-normal distributions [crow1987lognormal]. Note that we consider $(1-\gamma)$ to be sampled from a log-normal distribution, not $\gamma$, because observations (PSA levels before and after Abiraterone administration) suggest that $\gamma$ is typically much closer to $1$ than 0.
We then apply the LBEB approach [^3] We use the standard package nlmefit in MatLab R2023a [MATLAB] to algorithmically execute LBEB approach for NLME. to estimate the common parameters $\{r_{S}/r_{R},\alpha_{RS}\}$, the log-normal distributions of $\{r_{R},(1-\gamma),x_{S}(0),x_{R}(0)\}$, and the values $\{r_{R},\gamma,x_{S}(0),x_{R}(0)\}$ of each patient, and also the two constants $A$, $B$ that characterized our noise function, so that the observations across patients can be the most likely ones (maximum likelihood).
@@ IV.3 Best-Fit Results
We start by acknowledging that the clinical data is extremely noisy, which should be expected. The fit found in Zhang et al. [zhang2022evolution], while not quantitative, does capture nearly all the qualitative behavior of every patients regarding PSA response on- and off-drug. However, the total SSE for all patients of this fit is very large, which motivates us to reanalyze their data. While we still consider two common parameters, we now use four patient-specific parameters instead of only two, resulting in a significantly better fit, as shown in Fig. 2 and Fig. 3. The total SSE has been reduced by more than a factor of three, from $170.68$ for Zhang et al. [zhang2022evolution] fit to $57.52$ for our LBEB approach. The SSE for individual patients is also typically smaller using our approach (26 out of 32).
@@
Figure 2: Time series of PSA levels for the first 16 patients.. Each plot title shows the patient’s name, SSE for Zhang et al. [zhang2022evolution] fit, and the LBEB fit. The cyan line marks drug use periods, orange dots represent normalized PSA Eq. (10), the light-blue dashed line shows Zhang et al. [zhang2022evolution] best fit, and the blue line shows the LBEB best fit.
@@
Figure 3: Time series of PSA levels for the next 16 patients. Each plot title shows the patient’s name, SSE for Zhang et al. [zhang2022evolution] fit, and the LBEB fit. The cyan line marks drug use periods, orange dots represent normalized PSA Eq. (10), the light-blue dashed line shows Zhang et al. [zhang2022evolution] best fit, and the blue line shows the LBEB best fit.
@@
Figure 4: The relations between patient-specific parameters and the PSA level at the first drug treatment. Here we show the identified resistant cell growth rate $r_{R}$, drug effectiveness $\gamma$, and initial population values $\{x_{S}(0),x_{R}(0)\}$, and the PSA level when the total cancel population reaches its carrying capacity $\text{PSA}_{K}=\tilde{\lambda}$ for all patients. (A) The resistant cell growth rate $r_{R}$ and the drug ineffectiveness $1-\gamma$. The green continuous line shows a slight positive trend between these parameters in log-log plot ($R^{2}=0.10$ [montgomery2021introduction]), indicating a power-law relation $1-\gamma\propto r_{R}^{0.5\pm 0.5}$. (B) The initial values $\{x_{S}(0),x_{R}(0)\}$. The yellow dash line divides this two-dimensional parameter space into two: sensitive cell dominated (above) and resistant cell dominated (below), based on their influences on the growth of resistant cells at $t=0$. (C) The PSA levels at first drug treatment $\text{PSA}(t_{m^{\text{treat}}})$ and at carrying capacity $\text{PSA}_{K}$. The green continuous line shows a linear regression (that passes through the origin) between these parameters ($R^{2}=0.59$ [montgomery2021introduction]), indicating a linear relation $\text{PSA}(t_{m^{\text{treat}}})=\left(0.25\pm 0.05\right)\text{PSA}_{K}$. The yellow dash line divides this two-dimensional space into two: $\text{PSA}(t_{m^{\text{treat}}})/\text{PSA}_{K}>1/\alpha_{RS}$ (above) and $\text{PSA}(t_{m^{\text{treat}}})/\text{PSA}_{K}<1/\alpha_{RS}$ (below).
@@
We find the best common parameter values to be $r_{S}/r_{R}=2.97\pm 0.02$ and $\alpha_{RS}=5.2\pm 0.9$. This provides strong evidence that $r_{S}>r_{R}$. The large value for the competition coefficient $\alpha_{RS}$ between resistant and sensitive cells agrees with the findings of [zhang2022evolution], indicating the robustness of this estimation even with different approaches. The patient-specific parameters are sampled from
@@ IV.4 Comparison between Models
For linear regression, the standard metric for “goodness-of-fit” is given by the $R^{2}$-value [montgomery2021introduction]. A direct generalization of this dimensionless metric for nonlinear models [efron1978regression, anderson1994model] is commonly referred to as the Efron’s pseudo-$R^{2}$-value [schabenberger2001contemporary], is given by $\tilde{R}^{2}=1-\tilde{A}$. The coefficient of alienation $\tilde{A}$, defined as [anderson1994model]:
@@
where $\text{SS}_{\text{data}}$ is the data variance (i.e. SSE with respect to the mean value). This measures the proportion of residual variation in the nonlinear model compared to a null model with no explanatory power, indicating how much the model deviates from an ideal fit. An ideal fit corresponds to $\tilde{R}^{2}=1$. While there are many other different proposals for goodness-of-fit depending on the specific characteristics of the nonlinear model being used [mcfadden1972conditional, smith2013comparison, spiess2010evaluation], the Efron’s pseudo-$\tilde{R^{2}}-$value remains a useful metric due to its simplicity and ease of interpretation [anderson1994model].
Since the number of parameters in the fit models that we use are not the same, i.e. $n_{params}=(2\times N)+2=66$ in Zhang et al. [zhang2022evolution] and $n_{params}=(4\times N)+2=130$ in our approach, we need to calculate the adjusted Efron’s psuedo-$R^{2}$-value [archontoulis2015nonlinear] instead:
@@
We have $n_{\text{data}}=678$ and $\text{SS}_{\text{data}}=176.16$. The closer $\tilde{R}^{2}_{adj}$ is to $1$, the more we can trust a fitted model. We obtain $\tilde{R}^{2}_{adj}\approx-0.08$ for Zhang et al. [zhang2022evolution] fit and $\tilde{R}^{2}_{adj}\approx 0.60$ for our LBEB fit, which corroborates our approach.
@@ V Optimization of Treatment with Abiraterone
Within our adaptive therapy policy framework, our goal is to determine the optimal policy parameters $(\eta_{\uparrow},\eta_{\downarrow})$ that maximize the total achievable time $\tau_{\text{tot}}$ we can keep the cancer progression (with treatment) below the biomarker level $\text{PSA}_{\text{thr}}$, the level at which the patient begins to show symptoms, in order to avoid further complications. Consider the patients with two common parameters $\Theta=\{r_{S}/r_{R},\alpha_{RS}\}$ and the five patient-specific parameters $\{\text{PSA}_{K},\theta\}=\{\tilde{\lambda},r_{R},\gamma,x_{S}(0),x_{R}(0)\}$, as found according to Section IV.3. Note that $\text{PSA}_{K}=\tilde{\lambda}$ follows, as explained in Eq. (6). For simplification, we further assume that the PSA level is measured continuously, and that there is no noise in these measurements, so that the decision to turn on and off drug treatment based on the policy parameters as described in Eq. (8) is exact. Hence, we can calculate the total achievable time $\tau_{\text{tot}}$ below $\text{PSA}_{\text{thr}}$ using policy parameters $(\eta_{\uparrow},\eta_{\downarrow})$ for a patient $\mu$ with parameters $\{{\Theta,\text{PSA}_{K}}_{\mu},\theta_{\mu}\}$, [^4] We utilize the standard integrator ode23 in MatLab R2023a [MATLAB] to simulate Eq. (4) and implement the terminal condition, obtaining the total achievable time $\tau_{\text{tot}}$ at when the simulated PSA level crosses the threshold value $\text{PSA}_{\text{thr}}$. and we denote its full functional form as $\tau_{\text{tot}}^{\{{\Theta,\text{PSA}_{K}}_{\mu},\theta_{\mu}\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$. We can simplify this expression to $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$, and further drop all superscripts when it is clear which patient and what threshold are referred to.